Bernoulli's effect to measure speed of wind.Expain how it measures increases in speed of wind
1
Expert's answer
2020-02-24T10:51:48-0500

Bernoulli's principle states that an increase in the speed of a fluid occurs simultaneously with a decrease in pressure and vice versa. Specifically, if the speed of a fluid decreases to zero, then the pressure of the fluid will increase to its maximum. This is known as the stagnation pressure or total pressure. One special form of Bernoulli's equation is as follows:

Stagnation pressure = static pressure + dynamic pressure

"P_O=P_S+P_d=P_S+\dfrac{1}{2}\rho V^2"

where the stagnation pressure, Po, is the pressure if the flow speed is reduced to zero isentropically, the static pressure, PS, is the pressure the surrounding fluid is exerting on a given point, and the dynamic pressure, Pd, also called ram pressure, is directly related to the fluid density, ρ, and flow speed, V, for a given point. This equation only applies to incompressible flow, such as liquid flow and low-speed air flow (generally less than 100 m/s).

From the above equation, we can express flow speed, V, in terms of pressure differential and fluid density as:

"V=\sqrt{\dfrac{2(P_o-P_s)}{\rho}}"